OpenStudy (anonymous):

how do you solve log(10) 19

OpenStudy (anonymous):

Help @angela210793

OpenStudy (anonymous):

lets goo (:

OpenStudy (anonymous):

then our hero comes pls @Hero

hero (hero):

Hint: \[\log_{10}(19) = y \]

OpenStudy (anonymous):

are we tryin to find out y ??but hoow?

OpenStudy (anonymous):

i really need help with this

OpenStudy (anonymous):

what are trying to find out??

hero (hero):

Here's the formula: \[\log_x(a) = y \space \text{is equivalent to} \space x^y = a\]

hero (hero):

In other words, \[\log_{10} (19) = y \space \text {is equivalent to} \space 10^y = 19\]

hero (hero):

So basically we can just deal with \[10^y = 19\] Now use properties of logs to take y out of the exponent. You do that by taking logs of both sides: \[y \log(10) = \log(19)\] Then you isolate y: \[y = \frac{\log(19)}{\log(10)}\]

hero (hero):

You end up with y = 1.27875

hero (hero):

approximately